# Michel F. C. Haddad

### v1

@@ A.2 Sketch of proofs of asymptotic theories
In the panel case, this bias can be shown to depend on the difference between the conditional density and its kernel approximation:
$\rho_{d}(D_{t-1},X)-\rho_{d}^{h}(D_{t-1},X).$
Moreover, we can demonstrate that
$\rho_{d}^{h}(D_{t-1},X)$ $=E[\omega(D;d,h)|\mathbf{D_{t-1}},X]$
[…]
$=\rho_{d}(D_{t-1},X)+Ch^{2}\rho_{d}^{(2)}(D_{t-1},X)+o(h^{2}).$
For this reason, taking the difference $\rho_{d}(D_{t-1},X)-\rho_{d}^{h}(D_{t-1},X)$ and making use of the assumptions on the kernel function and the conditional density, the bias is of order $O(h^{2})$.
Secondly, to show asymptotic normality, we may employ the following decomposition:
@@
It suffices to focus on the difference between our ATET estimator and the kernel counterpart of the population ATET. By the definition of the score function (22) in Section 4, an application of Taylor’s theorem yields the following decomposition:
$\sqrt{n}(\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}-\Delta_{d_{t},d_{t}^{\prime},t}^{h})$ $=\sqrt{n}\sum_{i=1}^{n}E_{n}[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\hat{\eta})]$ (A.7)
$+\sqrt{n}\sum_{i=1}^{n}E_{n}[\partial_{P}\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\hat{\eta})](\hat{P}_{d_{t}}-P_{d_{t}})$ (A.8)
$+\sqrt{n}\sum_{i=1}^{n}E_{n}[\partial_{P}^{2}\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\bar{P},\hat{\eta})](\hat{P}_{d_{t}}-P_{d_{t}})^{2}$ (A.9)
where the derivative $\partial_{P}$ is w.r.t. the density $P_{d_{t}}$, and $\bar{P}\in(\hat{P}_{d_{t}},P_{d_{t}})$. We use $E_{n}$ to denote the empirical average. Furthermore, we omit a discussion of cross-fitting since it does not affect our main argument. Then, the asymptotic result follows from standard arguments and we only provide a high-level summary. In the first step, we show that the second-order term (A.9) vanishes asymptotically and does not contribute to the asymptotic variance. In the second step, we bound the first-order term (A.8) and show that this term contributes to the asymptotic variance. In the third step, we expand (A.7) around the nuisance estimator $\hat{\eta}$ and use Neyman orthogonality to show that the first-order bias disappears and second-order terms are negligible. In the final step, we employ a triangular array central limit theorem to establish asymptotic normality.
To establish the consistency of our variance estimator, we define:
$\tilde{\phi}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}})=\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)-\frac{\Delta_{d_{t},d_{t}^{\prime},t}^{h}}{P_{d_{t}}}\left(\omega(D_{t};d_{t},h)-E[\omega(D_{t};d_{t},h)]\right).$
@@
$\sigma_{h}^{2}=E[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)];\quad\hat{\sigma}_{h}^{2}=E_{n}[\tilde{\phi}^{2}(W,\hat{\Delta}_{d_{t},d_{t}^{\prime},t},\hat{P}_{d_{t}},\hat{\eta})].$
It has to be shown that
$J_{n}=|E_{n}[\tilde{\phi}^{2}(W,\hat{\Delta}_{d_{t},d_{t}^{\prime},t},\hat{P}_{d_{t}},\hat{\eta})]-E[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)]|=o_{p}(1).$
@@
$J_{n}$ $\leq|E_{n}[\tilde{\phi}^{2}(W,\hat{\Delta}_{d_{t},d_{t}^{\prime},t},\hat{P}_{d_{t}},\hat{\eta})]-E_{n}[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)]|$ (A.10)
$+|E_{n}[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)]-E[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)]|.$ (A.11)
We can demonstrate that (A.11) is $o_{p}(1)$ by making use of the assumptions on the kernel function and boundedness. Furthermore, one has to bound the term (A.10). We henceforth denote (A.10) by $J_{1n}$, which satisfies:
$J_{1n}^{2}\lesssim S_{n}\left(S_{n}+\frac{1}{n}\sum_{i=1}^{n}\tilde{\phi}^{2}(W_{i},\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)\right),$
@@
$S_{n}=\frac{1}{n}\sum_{i=1}^{n}|\tilde{\phi}(W_{i},\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)-\tilde{\phi}(W,\hat{\Delta}_{d_{t},d_{t}^{\prime},t},\hat{P}_{d_{t}},\hat{\eta})|^{2}.$
Specifically, $S_{n}$ may be of order $O_{p}(h^{-1}\varepsilon_{n}^{2}+h^{-2}n^{-1})$, which follows from an expension of the term $\tilde{\phi}(W,\hat{\Delta}_{d_{t},d_{t}^{\prime},t},\hat{P}_{d_{t}},\hat{\eta})$ around $\hat{\Delta}_{d_{t},d_{t}^{\prime},t}$ using Taylor’s theorem. Moreover, the remaining term in $J_{1n}^{2}$, namely $\frac{1}{n}\sum_{i=1}^{n}\tilde{\phi}^{2}(W_{i},\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)$, may be of order $O_{p}(h^{-1})$ by our assumptions on the kernel function and boundedness. Therefore, $J_{1n}$ is $o_{p}(1)$ if $h^{-2}\varepsilon_{n}^{2}+h^{-3}n^{-1}=o(1)$, and the consistency of our variance estimator follows.

### v2

@@ A.2 Proofs of asymptotic theories
In the repeated cross-sectional case, this bias can be shown to depend on the difference between the conditional density and its kernel approximation:
$\rho_{d_{t}}(D_{t-1},X)-\rho_{d_{t}}^{h}(D_{t-1},X).$
In particular, it can be shown that
$\rho_{d_{t}}^{h}(D_{t-1},X)$ $=E[\omega(D_{t};d_{t},h)\cdot I\{T=t\}|\mathbf{D}_{t-1},X]$
$=P(T=t|\mathbf{D}_{t-1},X)\int\frac{1}{h}\mathcal{K}\left(\frac{D_{t}-d_{t}}{h}\right)\rho_{s}(D_{t-1},T=t,X)ds$
$\asymp\rho_{d}(D_{t-1},t,X)+Ch^{2}\rho_{d}^{(2)}(D_{t-1},t,X)+o(h^{2}).$
Therefore, taking the difference $\rho_{d}(D_{t-1},t,X)-\rho_{d}^{h}(D_{t-1},t,X)$, by the assumptions on the kernel function and the conditional density, the bias is of order $O(h^{2})$.
Second, to establish asymptotic normality in Theorem 1, we consider the following decomposition:
@@
It suffices to focus on the difference between our ATET estimator and the kernel counterpart of the population ATET. In particular, the Neyman orthogonality results By the definition of the score function (22) in Section 5 and using Taylor’s theorem, we have the following expansion:
$\sqrt{n}(\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}-\Delta_{d_{t},d_{t}^{\prime},t}^{h})$ $=\sqrt{n}\frac{1}{K}\sum_{k=1}^{K}E_{nk}[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\hat{\eta}_{k})]$ (A.7)
$+\sqrt{n}\frac{1}{K}\sum_{k=1}^{K}E_{nk}[\partial_{\Pi}\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\hat{\eta}_{k})](\hat{\Pi}_{d_{t},t}^{k}-\Pi_{d_{t},t})$ (A.8)
$+\sqrt{n}\frac{1}{K}\sum_{k=1}^{K}E_{nk}[\partial_{\Pi}^{2}\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\bar{\Pi},\hat{\eta}_{k})](\hat{\Pi}_{d_{t},t}^{k}-\Pi_{d_{t},t})^{2}$ (A.9)
where the derivative $\partial_{\Pi}$ is w.r.t. the density $\Pi_{d_{t},t}$, and $\bar{\Pi}\in(\hat{\Pi}_{d_{t},t}^{k},\Pi_{d_{t},t})$. We use $E_{nk}$ to denote the empirical average with subsample $\mathcal{W}_{k}$. Moreover, due to cross-fitting, we can treat the high-dimensional nuisance parameters $\hat{\eta}_{k}$ as fixed without loss of generality. This significantly lessens the requirement on the complexity of the nuisance parameter space and greatly simplifies the asymptotic analysis. Then, the asymptotic result follows from standard arguments and we only provide a high-level summary. that by the standard arguments on kernel density, we have $|\hat{\Pi}_{d_{t},t}^{k}-\Pi_{d_{t},t}|=O_{p}((nh)^{-1/2})$ and $|\hat{\Pi}_{d_{t},t}^{k}-\Pi_{d_{t},t}|^{2}=O_{p}((nh)^{-1})$.
In the first step, we can show that the second-order term (A.9) vanishes asymptotically and does not contribute to the asymptotic variance. Specifically, by the standard results for kernel density estimators, $|\hat{\Pi}_{d_{t},t}^{k}-\Pi_{d_{t},t}|^{2}=O_{p}((nh)^{-1})$. Moreover, it can be shown that
$|E_{nk}[\partial_{\Pi}^{2}\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\bar{\Pi},\hat{\eta}_{k})]-E[\partial_{\Pi}^{2}\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta_{0})]|=O_{p}((nh)^{-1/2}+h^{-1/2}\varepsilon_{n}).$ (A.10)
Then by the boundness and rate assumptions, we conclude that (A.9) is $o_{p}(1)$.
In the second step, we bound the first-order term (A.8) and show that this term contributes to the asymptotic variance. It can be shown that
$E_{nk}[\partial_{\Pi}\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\hat{\eta}_{k})]=E[\partial_{\Pi}\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta_{0})]+o_{p}(1)$ (A.11)
if $h^{-1}\varepsilon_{n}^{2}=o(1)$. Then using the facts that $|\hat{\Pi}_{d_{t},t}^{k}-\Pi_{d_{t},t}|=O_{p}((nh)^{-1/2})$ and
$\hat{\Pi}_{d_{t},t}^{k}-\Pi_{d_{t},t}=\frac{1}{n-n_{k}}\sum_{i\in\mathcal{W}_{k}^{c}}\omega(D_{i,t};d_{t},h)\cdot I\{T_{i}=t\}-E[\omega(D_{i,t};d_{t},h)\cdot I\{T_{i}=t\}+O(h^{2}),$ (A.12)
we conclude that
[…]
where we have used the fact that $E[\partial_{\Pi}\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta_{0})]=\Delta_{d_{t},d_{t}^{\prime},t}^{h}/\Pi_{d_{t},t}$.
In the third step, we expand (A.7) around the nuisance estimator $\hat{\eta}$ and use Neyman orthogonality to show that the first-order bias disappears and second-order terms are negligible. Specifically, (A.7) can be decomposed as the following
[…]
where
$R_{nk}=E_{nk}[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\hat{\eta}_{k})]-E_{nk}[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta_{0})].$ (A.15)
To bound $R_{nk}$, we can apply triangle inequality, with $n_{k}:=n/K$ denoting the size of the subsample,
$|R_{nk}|\leq(R_{1k}+R_{2k})/\sqrt{n_{k}}$
where
$R_{1k}$ $=|G_{nk}[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\hat{\eta}_{k})]-G_{nk}[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta_{0})]|$ (A.16)
$R_{2k}$ $=\sqrt{n_{k}}|E[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\hat{\eta}_{k})|\mathcal{W}_{k}^{c}]-E[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta_{0})]|.$ (A.17)
Specifically, it can be shown that $R_{1k}=O_{p}(h^{-1}\varepsilon_{n})$ by sample splitting, the assumptions on the nuisance estimators, the law of iterated expectation, and conditional Markov’s inequality. Moreover, by Neyman orthogonality, it can be shown that $R_{2k}=O(n^{1/2}h^{-1/2}\varepsilon_{n}^{2})$. Therefore,
[…]
In the final step, combining above results, the asymptotic normality follows from a triangular array central limit theorem.
Lastly, to establish the consistency of our variance estimator as in Theorem 2, we define:
$\tilde{\phi}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)=\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)-\frac{\Delta_{d_{t},d_{t}^{\prime},t}^{h}}{\Pi_{d_{t},t}}\left(\omega(D_{t};d_{t},h)-E[\omega(D_{t};d_{t},h)]\right).$
@@
[…]
It suffices to show that
$J_{n}:=|E_{n}[\tilde{\phi}^{2}(W,\hat{\Delta}_{d_{t},d_{t}^{\prime},t},\hat{\Pi}_{d_{t},t},\hat{\eta})]-E[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)]|=o_{p}(1).$
@@
$J_{n}$ $\leq|E_{n}[\tilde{\phi}^{2}(W,\hat{\Delta}_{d_{t},d_{t}^{\prime},t},\hat{\Pi}_{d_{t},t},\hat{\eta})]-E_{n}[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)]|$ (A.19)
$+|E_{n}[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)]-E[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)]|.$ (A.20)
First, we can show that (A.20) is $o_{p}(1)$ using the assumptions on the kernel function and boundedness. Moreover, the term (A.19), which we denote by $J_{1n}$, satisfies:
$J_{1n}^{2}\lesssim S_{n}\left(S_{n}+\frac{1}{n}\sum_{i=1}^{n}\tilde{\phi}^{2}(W_{i},\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)\right),$
@@
$S_{n}=\frac{1}{n}\sum_{i=1}^{n}|\tilde{\phi}(W_{i},\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)-\tilde{\phi}(W,\hat{\Delta}_{d_{t},d_{t}^{\prime},t},\hat{\Pi}_{d_{t},t},\hat{\eta})|^{2}.$
[…]
A.3 Descriptive statistics
Table 4: Descriptive statistics for extended set of variables
variable mean median std. dev. skewness kurtosis minimum maximum
Panel A: COVID-19 outcomes
deaths / population 0.0020 0.0019 0.0014 0.50 -0.62 0.0000 0.0075
Panel B: vaccination
dose-1 vaccination 0.26 0.11 0.30 0.79 -0.98 0.00 1.30
dose-2 vaccination 0.36 0.24 0.35 0.27 -1.65 0.00 1.07
Panel C: Human Development Index (MHDI)
municipal HDI (overall) 0.74 0.75 0.05 -0.69 0.77 0.57 0.86
longevity dimension 0.84 0.84 0.03 -0.74 0.62 0.72 0.89
education dimension 0.67 0.68 0.07 -0.61 0.45 0.42 0.81
income dimension 0.73 0.74 0.06 -0.34 0.73 0.56 0.89
Panel D: inequality & poverty
Gini index 0.50 0.50 0.05 -0.09 -0.15 0.35 0.63
Theil-L index 0.46 0.45 0.10 0.34 -0.01 0.21 0.76
proportion of poor 8.69 5.48 8.08 1.83 3.25 0.44 43.46
proportion extremely poor 2.72 1.29 3.67 2.82 10.08 0.00 25.72
vulnerable to poverty (%) 24.00 19.88 14.17 1.10 0.60 3.85 69.10
ratio 10% rich / 40% poor 13.58 12.79 4.10 0.92 1.03 5.98 28.23
per capita income 807.73 792.54 282.31 1.12 2.75 257.64 2043.74
income approp. (Top 10%) 40.75 40.70 4.86 0.06 -0.23 28.77 53.15
income approp. (Bottom 40%) 12.65 12.61 2.23 0.14 -0.08 6.98 19.26
Panel E: education & literacy (%)
complete secondary (25+) 36.20 36.22 8.89 0.14 0.17 15.59 64.08
complete primary (25+) 51.78 52.44 9.90 -0.30 0.00 23.79 76.83
complete higher ed (25+) 11.16 10.59 5.05 0.99 1.91 1.57 31.86
illiteracy rate (11-14) 1.90 1.34 1.68 3.19 13.20 0.00 13.04
illiteracy rate (15+) 7.26 5.65 5.07 2.15 5.14 1.55 29.69
Panel F: school attendance rates (%)
age 0 to 3 27.66 26.28 10.20 0.45 -0.17 7.31 56.87
age 15 to 17 84.36 85.13 4.35 -0.56 0.13 72.25 95.56
age 18 to 24 29.68 30.16 5.94 -0.09 -0.10 12.81 45.85
age 4 to 5 83.40 85.02 10.34 -0.96 0.76 43.36 98.88
age 6 to 14 97.51 97.61 1.09 -0.93 2.15 92.37 99.76
net enrollment primary 92.57 92.61 1.77 -0.10 0.46 86.52 97.76
net enrollment secondary 49.03 49.53 9.31 -0.18 -0.48 23.94 72.90
net enrollment higher ed 15.85 15.53 6.42 0.37 0.07 2.74 35.89
Panel G: demographics & survival
total population 2022 211,866 103,440 306,890 4.11 25.44 2,892 2,817,381
urban population (prop.) 0.90 0.93 0.10 -1.94 5.18 0.36 1.00
rural Population (prop.) 0.10 0.07 0.10 1.94 5.18 0.00 0.64
aging rate 11.46 11.67 2.71 -0.22 -0.11 3.67 18.73
total fertility rate 1.81 1.79 0.30 1.07 2.46 1.25 3.07
dependency ratio 43.71 43.30 2.95 0.38 -0.33 37.64 52.13
prob. survival to 40 yrs 94.44 94.64 1.18 -0.68 0.04 91.11 96.66
prob. survival to 60 yrs 84.56 84.83 2.22 -0.36 0.13 77.37 89.79
life expectancy at birth 75.20 75.42 1.72 -0.74 0.62 68.36 78.44
under-five mortality rate 16.48 15.63 3.73 1.78 5.96 9.98 38.14
infant mortality rate 14.55 13.80 3.59 1.80 5.66 8.49 35.00
Panel H: housing & infrastructure (%)
households piped water 95.63 97.80 5.56 -2.98 12.26 60.18 99.99
households density >> 2 22.82 22.00 9.03 0.62 0.06 6.55 51.40
households electricity 99.72 99.90 0.50 -3.75 16.50 96.48 100.00
households garbage coll. 98.35 99.34 2.72 -3.59 17.34 78.10 100.00
households sewage 68.49 78.72 28.91 -0.93 -0.34 0.71 99.83
sewage / septic tank 84.60 90.51 17.77 -2.05 4.28 5.30 99.95
inadequate water/sewage 2.55 0.63 4.83 3.17 10.72 0.00 27.16
non-masonry walls 1.54 0.65 2.76 4.87 32.22 0.00 25.84